OpenStudy (steve816):

A force of F on a mass M causes an acceleration of 2.3 m/s/s. What acceleration will result from a force of two-thirds F on a mass of one-third M?

OpenStudy (steve816):

I know that F=m*2.3

OpenStudy (mhchen):

So isn't that \[\frac{2F}{3} = \frac{m}{3}*2.3\\2F = m*(2.3*3)\\F=m*(\frac{2.3*3}{2})\]

OpenStudy (mathmate):

Right, F=Ma, so a=F/M=2.3 => acceleration=(2/3)F/(M/3)=2F/M=?

OpenStudy (steve816):

Wait, so how do I solve for acceleration when there are two variables???

OpenStudy (mathmate):

You don't! The ratio of the variables is known (F/M). And acceleration is an expression in terms of the same ratio.

OpenStudy (steve816):

Ah, I got the right answer! Thank you mathmate!

OpenStudy (mathmate):

You're welcome! :)